# Conformal Tradeoffs: Guarantees Beyond Coverage

### v1

@@ H.4 Representation and model protocol
H.5 Conformal calibration and evaluation
@@ Appendix I Solubility supplementary details
This appendix records methodological details for the solubility scenario-planning experiments in Section 5.3. The purpose is reproducibility and documentation.
I.1 Operational outcome definitions
I.2 Pareto-front artifact and representative regimes
The full deduplicated Pareto front used in Section 5.3 is provided as a CSV artifact (pareto_full.csv). Each row corresponds to a distinct Pareto-optimal operating point after deduplication in KPI space. The CSV records (i) the SSBC effective levels $(\alpha_{\mathrm{SSBC},0},\alpha_{\mathrm{SSBC},1})$, (ii) the nominal request parameters $(\alpha_{0},\delta_{0},\alpha_{1},\delta_{1})$ that induced the operating point, (iii) the six joint outcome rates $P(Y=y,C=c)$, and (iv) predictive-envelope endpoints for each joint rate (LB/UB), reported on the same probability scale and then converted below to counts per 1000.
Table 9 shows two representative Pareto regimes drawn from the provided Pareto front to illustrate contrasting planning postures: a loss-minimizing regime (low irreversible exclusion of soluble compounds) and a decisiveness-maximizing regime (low total hedging). Counts are expected events per $1000$ molecules; brackets denote the corresponding predictive envelopes.
Table 9: Two representative Pareto-optimal operating regimes for the solubility scenario, illustrating contrasting planning postures. Rates are expected counts per 1000 molecules, with predictive envelopes in brackets (derived from pareto_full.csv).
Loss-minimizing regime High-decisiveness regime
$\alpha_{0}$ 0.125 0.150
$\delta_{0}$ 0.150 0.150
$\alpha_{1}$ 0.050 0.150
$\delta_{1}$ 0.075 0.100
Rate Interval Rate Interval
Loss rate 3 [0, 35] 12 [0, 58]
Waste rate 70 [27, 130] 13 [0, 49]
Total hedge rate 785 [663, 896] 514 [376, 675]
Correct soluble singleton rate 7 [0, 38] 88 [40, 155]
Correct insoluble singleton rate 135 [68, 215] 282 [196, 375]
Decisiveness ($1-\text{hedge}$) 215 486
I.3 Dataset and label construction
@@
I.4 Partitioning: scaffold-isolated training and exchangeable evaluation
We use a two-stage split to separate generalization across chemical space from the exchangeable design required for conformal calibration and finite-window operational analysis:
  • Scaffold grouping. Molecules are grouped by Bemis–Murcko scaffolds (Bemis and Murcko, 1996) computed from RDKit (Landrum, 2025) SMILES.
  • Scaffold-isolated training set. Scaffold groups are randomly shuffled and assigned to the training set until approximately 70% of molecules are allocated, ensuring no scaffold in training appears in the remaining pool.
  • Exchangeable calibration/test sets. The remaining scaffold-disjoint molecules are split into calibration and test sets by a stratified random split (50/50 of the remaining pool), preserving class proportions. These two sets are treated as exchangeable with respect to each other and to future draws under the evaluation regime in Section 5.3.
I.5 Molecular representation
@@
I.6 Model and training protocol
We train a CatBoost (Prokhorenkova et al., 2018) gradient-boosted decision tree classifier on the scaffold-isolated training set with fixed hyperparameters (depth 6, learning rate 0.05, multi-class loss), up to 1000 boosting iterations with early stopping (patience 50) using calibration-set performance. After training, the predictor is frozen and treated as infrastructure; the experiments study the conformal and operational layers rather than optimizing base accuracy.
I.7 Calibration and operationalization
Uncertainty quantification uses split conformal prediction with SSBC calibration (Section 5.1.1 and Appendix C), which selects a grid index to stabilize realized coverage semantics at user-specified confidence (Vovk, 2012b; Yang and Kuchibhotla, 2021). This calibration layer supports the finite-window predictive-envelope constructions used for operational planning.
@@ I.7 Calibration and operationalization
I.8 Deployment-matched calibration via chemical tribes
@@
Generic scaffold attribution.
For interpretability we compute a generic scaffold representation by extracting the Murcko scaffold and converting it to a generic form (all atoms carbon, all bonds single). This increases scaffold group size and supports qualitative attribution across broader structural families.

### v2

@@ H.3 Representation and model protocol
H.4 Conformal calibration and evaluation
@@ Appendix I Solubility supplementary details
This appendix records methodological details for the solubility scenario-planning experiments in Section 5.3. We first describe the dataset, partitioning, modeling, and calibration setup, and then document the operational artifacts used to interpret the resulting Pareto front.
I.1 Dataset and label construction
@@
I.2 Partitioning: random training and exchangeable evaluation
We use a random split to separate model fitting from the exchangeable design required for conformal calibration and finite-window operational analysis:
  • Random train/calibration/test split. Molecules are partitioned by a stratified random split, with approximately 70% assigned to training and the remaining molecules split equally between calibration and test, preserving class proportions.
  • Exchangeable calibration/test sets. The calibration and test sets are treated as exchangeable with respect to each other and to future draws under the evaluation regime in Section 5.3.
I.3 Molecular representation
@@
I.4 Model and training protocol
We train a CatBoost (Prokhorenkova et al., 2018) gradient-boosted decision tree classifier on the training set with fixed hyperparameters (depth 6, learning rate 0.05, multi-class loss), up to 1000 boosting iterations with early stopping (patience 50) based on a separate hyperparameter tuning-set performance metric. After training, the predictor is frozen and treated as infrastructure; the experiments study the conformal and operational layers rather than optimizing base accuracy.
I.5 Calibration and operationalization
Uncertainty quantification uses split conformal prediction with SSBC calibration (Section 5.1.1 and Appendix C), which selects a grid index to stabilize realized coverage semantics at user-specified confidence (Vovk, 2012b). This calibration layer supports the finite-window predictive-envelope constructions used for operational planning.
@@ I.5 Calibration and operationalization
I.6 Deployment-matched calibration via chemical tribes
@@
I.7 Operational outcome definitions
I.8 Pareto-front artifact and representative regimes
With the methodological setup above fixed, Section 5.3 produces a deduplicated Pareto front in KPI space by sweeping SSBC calibration settings and retaining nondominated operating points. Table 9 shows two representative Pareto regimes drawn from that front to illustrate contrasting planning postures: a loss-minimizing regime (low irreversible exclusion of soluble compounds) and a decisiveness-maximizing regime (low total hedging). Counts are expected events per $1000$ molecules; brackets denote the corresponding predictive envelopes.
Table 9: Two representative Pareto-optimal operating regimes for the solubility scenario, illustrating contrasting planning postures. Rates are expected counts per 1000 molecules, with predictive envelopes in brackets.
Loss-minimizing regime High-decisiveness regime
$\alpha_{0}$ 0.125 0.150
$\delta_{0}$ 0.150 0.150
$\alpha_{1}$ 0.050 0.150
$\delta_{1}$ 0.075 0.100
Rate Interval Rate Interval
Loss rate 3 [0, 35] 12 [0, 58]
Waste rate 70 [27, 130] 13 [0, 49]
Total hedge rate 785 [663, 896] 514 [376, 675]
Correct soluble singleton rate 7 [0, 38] 88 [40, 155]
Correct insoluble singleton rate 135 [68, 215] 282 [196, 375]
Decisiveness ($1-\text{hedge}$) 215 486